OpenStudy (anonymous):

The measure of two angles of a triangle are 35 degrees and 45 degrees. What is the measure of the third angle of the triangle? How to solve

OpenStudy (ivettef365):

what is the total measure of the angles of a triangle ???

OpenStudy (anonymous):

Every triangle has 180 degrees on the inside, so you take a+b+c=180. Now you already have two angles so you plug them in: 35+45+c=180. Now you can add like terms on the left side: 80+c=180 Now you can subtract 80 from the left side to the right side. 80-80+c=180-80 c=100 degrees.

OpenStudy (anonymous):

Thank you!